# Is Long Context All You Need? Leveraging LLM’s Extended Context for NL2SQL Experiments, Analysis and Benchmark Paper

### v1

@@ 3.2. Fix & Rewrite
We allow the model to fix its errors and rewrite the SQL output based on execution. If the generated SQL query returns an error during execution, it triggers correction modules (Wang et al., 2024; Caferoğlu and Ulusoy, 2024) and retries until a valid SQL is returned for a fixed number of times . (Pourreza and Rafiei, 2024) showed that LLMs can correct for minor syntax errors in SQL. While the existing self-correction mechanisms and the latest LLMs themselves are effective in fixing SQL syntax errors, they do not address more nuanced semantic errors like invalid literal references and incorrect join paths. Semantic errors may persist after syntax error correction (e.g., referencing a valid literal value for a wrong column) and they are hard to detect without the ground-truth query results. We check for potential semantic errors when the query returns empty result; we ask the model to rewrite the query given that the previous query resulted in an empty result set and with more extensive lists of sample column values. There is a risk of false positive in relying on null (an empty result) for detecting semantic errors. In a case that the ground truth should be null, this would incur additional cost of repeating the task. If the question is not ambiguous, the model should return the same correct null query even with the extra sample column values. If the question is ambiguous and the model re-generates a non-empty result, then we prefer that over the null and move to the next verification step. Seeing the full list of column values enables the model to select the right keyword and columns and to reason about more correct alternative join paths. Passing the full lists of sample column values inflate the context size significantly and requires long context and LLMs that can retrieve accurately over the long context. We evaluate the impact of this long context disambiguation strategy in Section 4.6. This technique is expensive, but it can complement imperfect table & column selection and schema linking, which often leads to sub-optimal NL2SQL performance (Caferoğlu and Ulusoy, 2024).
@@ 3.3. Verify
We use the untuned gemini-1.5-pro LLM to verify the correctness of the final output. The model is given the entire database table schema and question (with additional hints if provided) to judge (see Appendix Fig. 10 for the prompt). We believe that fine-tuning a verifier or a picker for self-consistency(Wang et al., 2023; Pourreza et al., 2024; Zheng et al., 2024) can further improve the accuracy. While we focus on leveraging the long context ICL for a typical NL2SQL pipeline with simpler techniques (self-correction, verification without fine-tuning), the aforementioned techniques are orthogonal and can be used in conjunction. We demonstrate how this verification step can impact the final performance and the generation cost in Section 5.
@@ 4.1. Evaluation setup
We use public Google Cloud (GCP) Vertex AI gemini API for all our experiments. This is important for reproducibility of our experiments and analysis. We use the latest publicly available gemini-1.5-pro-002 and gemini-1.5-flash-002 checkpoints. gemini-1.5 is a long-context LLM class with up to 2-million token context for pro and 1-million token context for flash. Our test VMs and Vertex AI endpoints are located in the same GCP region (us-central1-b).
We report the full pipeline performances on various NL2SQL benchmark datasets (BIRD (Li et al., 2024a), SPIDER 1.0 (Yu et al., 2018), KaggleDBQA (Lee et al., 2021) and BEAVER (Chen et al., 2024)) in Section 4.2. We run our micro-benchmark experiments using BIRD dev as a representative dataset, since it contains the most number of questions with varying degrees of difficulties spanning over multiple tables with varying schema complexity. BIRD is also currently the most popular NL2SQL benchamrk with a leaderboard (all the latest NL2SQL research publications use this dataset to compare their performances).
For the performance and cost metrics, we use popular execution accuracy (Ex Acc) and the accumulated tokens per request ($L$), and/or normalized latency per request ($T$) as the absolute latency measure behind the public API endpoints can vary from time to time – a single gemini-1.5-pro request latency with 8k-token is used as a reference unit ($T$=1.0); we report floating-point values to one decimal place, ignoring differences of less than 10%. Likewise, we refrain from reporting the monetary cost directly, which depends on the pricing model (Google Cloud, 2024a) that is subject to change. Instead, we report the accumulated number of tokens per request $L$, which is highly correlated to the total monetary cost and invariant to the pricing model change.
There is a strong positive correlation between $L$ and $T$, as discussed in Section 4.8, and we report both in the ablation study. We report the average metrics across requests with tight variance (margin of error/variability less than 2%); if the observed variance is large, we report the mean plus one standard deviation (e.g., $\bar{L}+\sigma$) to bound the majority of requests.
Our “full” pipeline employs all the information and the techniques as discussed in Section 3 and Section 5; for the individual experiments we use the “baseline” pipeline to study specific questions being addressed. The baseline pipeline includes entire DB schema, sample column values, instructions and hints, but excludes self-correction, disambiguation, synthetic examples and no verification.
@@ 4.2. Benchmark evaluation
Table 2. Benchmark performance and generation latency $T_{g}$ using the full long-context NL2SQL pipeline and with gemini-1.5 models. $T_{g}$ measures per request SQL output generation time, including time spent on self-correction and retries.
@@
SPIDER 1.0 test 85.70 1.5 84.60 1.1
KaggleDBQA test 61.10 2.0 59.50 1.6
BEAVER dw [^1] Only Data Warehouse 1 (dw) dataset with 48 questions is publicly available as of January 7th, 2025. 47.7 371.3 45.5 426.3
Table 2 shows the full pipeline evaluation results with various benchmark datasets. The popular benchmark datasets (BIRD dev, SPIDER 1.0 test, KaggleDBQA test) provide a good set of questions of mixed difficulties over multiple domains and put our performance in perspective with other state-of-the-arts. Our approach leveraging the extended context of gemini-1.5 model can yield very competitive results in all the benchmarks, without techniques like fine-tuning and self-consistency (Table 1). The newer benchmark for enterprise warehouse data (BEAVER dw) is also interesting because the business-oriented questions and real enterprise tables are more complex and larger – BIRD is the most complex among the three popular benchmarks, and on the average, its questions have 0.918 JOINs over 6.82 tables/DB; BEAVER questions have 4.25 JOINs over 105.0 tables/DB. BEAVER is a new benchmark and is still in early stage. The best 1-shot Ex Acc on BEAVER (dw + 45 additional questions) from (Chen et al., 2024) is 2.0 in PL/SQL using gpt-4o with relevant table schema retrieval and 0.0 without retrieval.
The normalized generation latency $T_{g}$ increases with both larger context size and task difficulty (question and schema complexity), as it includes the self-correction and retries (with exponential back-off and max retries per request set to 5). For instance, BEAVER takes much longer to generate on average as our long context pipeline produces context size exceeding the limits (2-million tokens for pro and 1-million tokens for flash) or results in semantically invalid (empty) outputs, in which case the model retries a fixed number of times until it finds a better answer. For this reason, flash can take longer on average than pro.
In general, gemini-1.5-pro is better at complex reasoning tasks like NL2SQL, whereas gemini-1.5-flash is more cost-effective (lower pricing per token). We have noticed that the latest releases of the gemini models (released for public access on Sep. 24th, 2024) show significant performance improvements, and most notably, flash now performs more comparably to pro on the select benchmarks. This makes the more cost-efficient flash a very attractive choice for production NL2SQL.
@@ 4.3. Schema linking with all database tables
We explore the impact of passing the entire table schema set collected from entire databases. Schema linking is such a critical component that we strive to achieve a high-recall in production. With a long context model, we can pass a lot more table definitions and schema more loosely than the most relevant top-k tables. As with a standard practice, we retrieve the most relevant tables based on question and full table creation statement with column values and descriptions, embedding similarity. All the benchmark datasets provide specific target DB domain per question, which is different from the production setup we assumed in Figure 2 for the illustration. This narrows the search space, but even so, getting the perfectly accurate retrieval results is often infeasible. The goal is to ensure near perfect (higher) recall at the cost of lower precision by providing more tables, if not all, via long context .
Table 3. Table retrieval (TBR) performance and BIRD dev execution accuracy with the baseline pipeline. Top-K table retrieval ($k$=1,7) was done across the specific target DB tables per question; putting all tables (all tables/DB) guarantees perfect recall
# tables (k) k=1 k=7 all DB tables all dev tables
@@
Context size (tok/req) 2002.54 4627.68 7380.86 72619.96
The results shown in Table 3 demonstrates how the model does not get confused with a large number of mostly irrelevant table definitions in the context; thus providing more tables rather than solely relying on inaccurate table retrieval mechanism could make sense if the context size permits. The result ($k=1$) also re-affirms that schema linking is critical, since without all the ground truth tables in the prompt ($k=1$) the model cannot generate good quality SQL outputs. On the flip side, providing all tables from the target DB ensures perfect TBR recall and results in higher Ex Acc. This means that TBR recall is more important than precision or accuracy as the model generation quality does not degrade much with many irrelevant tables (93.12 irrelevant tables on the average). all DB tables use up to 13 tables to each request depending on the target DB; the average number of tables across different DBs is 6.82. Notice that the execution accuracy result for all dev tables (100 tables from BIRD dev) is slightly higher (within 5% margin of error – we use medium temperature for LLM generation and Ex Acc vary slightly per run), which indicates that the results are comparable. However, the average context size increases significantly as with all dev tables.
On the one hand, it is better to include as many tables (e.g., user history, all other tables within the target DB) to guarantee high TBR recall, which should be possible with long-context LLMs. On the other hand, the latency cost of processing larger context is higher (see Fig. 5) – and to some extent, adding more irrelevant tables would not help and justify the increase cost. Furthermore, some production setting require processing hundreds of very wide tables, which prevents putting all the tables in the context prohibitively expensive (see Section 6.3 for more discussion). Thus, accurate retrieval/filtering is still desirable and one should consider leveraging long-context to offset the imperfect retrieval mechanism for higher recall.
@@ 4.4. Many-shot ICL with example selection
Here we evaluate the impact of many-shot in-context learning (ICL) using examples selected from BIRD train dataset. The examples are retrieved based on the question embedding similarity.
Table 4. BIRD dev Ex Acc (%) with the baseline pipeline & selected similar examples from BIRD train dataset, $\sigma$(train), and ground-truth (+ GT)
@@
Context size (tok/req) 7380.86 8000.67 9927.24 13808.71 20358.30
(b) GT inserted at different positions (normalized) within the context window with selected 100 BIRD train examples
@@
$\sigma$(train) + GT 77.77 78.29 78.1 78.42 78.16
(c) Changing the order of the examples block (100 BIRD train examples with GT) in the context: beginning (before the instructions), middle (before the schema details), end (after the schema details)
@@
$\sigma$(train) + GT 78.62 78.42 78.62
Table 4(c)(a) shows that the train examples did not provide much boost and instead resulted in worse performance. This means that the model’s ability to learn from similar examples from BIRD train is quite limited and can be distracted from them. Also notice that the impact of distraction seems to be limited, say going from 1 similar example (11k tokens) to 100 similar examples (50k tokens). To examine the model’s sensitivity to the many irrelevant train examples further, we injected 1 ground-truth example from BIRD dev in the “middle” of the example blob (the average context size is omitted as it should be almost identical to the previous table). While the model can select the ground truth example among many other irrelevant ones, the presence of those irrelevant train examples can confuse the model as opposed to just having that ground truth. The recall matters more than precision, but the model is still sensitive to bad examples (low precision) to a limited degree – note that the accuracy with 100 random examples and 1 ground truth is much higher than no examples at all.
It is also important to note that the model is robust to the position and ordering of the ground-truth example. (Liu et al., 2024a) showed that LLMs tend to put more emphasis on the information in the beginning and the end, where the order of information in context matters and calls proper ranking and ordering for RAG. As in Table 4(c)(b-c), where the GT is injected at different positions within the context window, we observe that with gemini-pro-1.5 the ordering of the examples (and the schema information relocated by the relocation of the examples in (c)) has a negligible impact, and the long-context LLMs can retrieve the relevant ground truth example (also the schema information) from anywhere within its context window.
@@
Figure 4. Question similarity-based example selection ($\sigma$) vs. online example generation. The baseline uses no examples; examples were retrieved from BIRD train dataset, dev dataset and synthetic generated examples. train + GT uses the retrieved train examples and the ground-truth SQL from dev.
Fig. 4 illustrates how synthetic examples improve the quality of generated SQL outputs and compare with other types of examples retrieved from different example pools. As with a standard practice, the example retrieval was done via question embedding similarity using gecko embedding model. Note that using the ground-truth SQL queries from dev dataset provides performance ceilings ($\sigma$(dev) and $\sigma$(train + GT)), but is not allowed without an oracle (i.e., passing the answer as part of the request prompt). The gap between $\sigma$(dev) and $\sigma$(train + GT) indicates that providing the GT as one of the many examples is better than retrieving relevant examples from the pool that contains the GT, due to the imperfect retrieval process. $\sigma$(train) is the common practice where the relevant examples are selected from available datasets, such as the training dataset (train). Providing relevant train examples does as least as good as the baseline (no example) with a slight improvement when fewer than 200 shots are provided. Interestingly, synthetic examples ($\sigma$(synthetic) and original synthetic) provide significant uplift over $\sigma$(train). The gap between $\sigma$(synthetic) and original synthetic indicates that one can be more economical with the context size (achieving higher performance with a fewer examples) by filtering the relevant synthetic examples.

### v3

@@ 3.2. Fix & Rewrite
We allow the model to fix its errors and rewrite the SQL output based on execution. If the generated SQL query returns an error during execution, it triggers correction modules (Wang et al., 2024; Caferoğlu and Ulusoy, 2024) and retries until a valid SQL is returned within a fixed number of times . (Pourreza and Rafiei, 2024) showed that LLMs can correct minor syntax errors in SQL. While the existing self-correction mechanisms and the latest LLMs themselves are effective in fixing SQL syntax errors, they do not address more nuanced semantic errors like invalid literal references and incorrect join paths. Semantic errors may persist after syntax error correction (e.g., referencing a valid literal value for a wrong column), and they are hard to detect without the ground-truth query results. We check for potential semantic errors if the query returns empty result; we ask the model to rewrite such queries by providing more extensive list of sample column values. There is a risk of false positives in relying on a empty results for detecting semantic errors. If the correct ground-truth result is empty, this could incur additional computational costs due to unnecessary retries. If the question is not ambiguous, the model should return the same correct empty result even when additional sample column values are included. However, if the question is ambiguous and the model re-generates a non-empty result, then we prefer that over empty result and move to the next verification step. Seeing the full list of column values enables the model to select the right keyword and columns and to reason about more accurate alternative join paths. Passing the full lists of sample column values inflate the context size significantly and requires long context and LLMs that can retrieve accurately over the long context. We evaluate the impact of this long context disambiguation strategy in Section 4.6. This technique is expensive, but it can complement imperfect table & column selection and schema linking, which often leads to sub-optimal NL2SQL performance (Caferoğlu and Ulusoy, 2024).
@@ 3.3. Verify
We use the untuned gemini-1.5-pro LLM to verify the correctness of the final output. The model is given the entire database table schema and question (with additional hints, if provided) to judge. We believe that fine-tuning a verifier or a picker for self-consistency(Wang et al., 2023; Pourreza et al., 2024; Zheng et al., 2024) can further improve accuracy. While we focus on leveraging the long context ICL for a typical NL2SQL pipeline with simpler techniques (self-correction, verification without fine-tuning), the aforementioned techniques are orthogonal and can be used in conjunction. We demonstrate how this verification step can impact the final performance and the generation cost in Section 5.
@@ 4.1. Evaluation setup
We use the public Google Cloud (GCP) Vertex AI gemini API for all our experiments. This is important for reproducibility of our experiments and analysis. We use the latest publicly available gemini-1.5-pro-002 and gemini-1.5-flash-002 checkpoints. gemini-1.5 is a long-context LLM class with up to 2-million token context for pro and 1-million token context for flash. Our test VMs and Vertex AI endpoints are located in the same GCP region (us-central1-b).
We report the full pipeline performance on various NL2SQL benchmark datasets (BIRD (Li et al., 2024a), SPIDER 1.0 (Yu et al., 2018), KaggleDBQA (Lee et al., 2021) and BEAVER (Chen et al., 2024)) in Section 4.2. We run our micro-benchmark experiments using two widely used benchmark datasets, BIRD dev and KaggleDBQA test. BIRD is currently the most popular NL2SQL benchmark with a leaderboard (all the recent NL2SQL research publications use this dataset to compare their performance), and it contains the largest number of questions with varying degrees of difficulties spanning over multiple tables with varying schema complexity. KaggleDBQA is another benchmark dataset for Text-to-SQL tasks that focuses on real-world, cross-domain data sources.
For the performance and cost metrics, we use popular execution accuracy (Ex Acc) and the accumulated tokens per request ($L$), and/or normalized latency per request ($T$) as the absolute latency measure behind the public API endpoints can vary from time to time – a single gemini-1.5-pro request latency with 8k-token is used as a reference unit ($T$=1.0); we report floating-point values to one decimal place, ignoring differences below 10%. Likewise, we refrain from reporting the monetary cost directly, as it depends on the pricing model (Google Cloud, 2024a), which is subject to change. Instead, we report the accumulated number of tokens per request $L$, which is strongly correlated with total monetary cost and remains invariant to pricing model changes.
There is a strong positive correlation between $L$ and $T$, as discussed in Section 4.8, and we report both in the ablation study. We report the average metrics across requests with low variance (margin of error/variability less than 2%); if the observed variance is large, we report the mean plus one standard deviation (e.g., $\bar{L}+\sigma$) to bound the majority of requests.
Our “full” pipeline employs all the contextual information and the techniques as discussed in Section 3 and Section 5; for the individual experiments, we use the “baseline” pipeline to study specific questions being addressed. The baseline pipeline includes entire DB schema, sample column values, instructions and hints, but excludes self-correction, disambiguation, synthetic examples and verification.
@@ 4.2. Benchmark evaluation
Table 2. Benchmark performance and generation latency $T$ using the full long-context NL2SQL pipeline and with gemini-1.5 models. $T_{g}$ measures per request SQL output generation time, including time spent on self-correction and retries.
@@
SPIDER 1.0 test 87.10 1.5 84.60 1.1
KaggleDBQA test 61.10 2.7 58.40 1.8
BEAVER dw [^1] Only Data Warehouse 1 (dw) dataset with 48 questions is publicly available as of January 7th, 2025. 60.41 506.1 50.00 506.1
Table 2 shows the full pipeline evaluation results across various benchmark datasets. The popular benchmark datasets (BIRD dev, SPIDER 1.0 test, KaggleDBQA test) provide a diverse set of questions with varying difficulties across multiple domains, allowing us to compare our performance with other state-of-the-art approaches. Our approach leveraging the extended context of gemini-1.5 model can yield competitive results across all the benchmarks, without techniques like fine-tuning and self-consistency (Table 1). The newer benchmark for enterprise warehouse data (BEAVER dw) is particularly noteworthy because its business-oriented questions and real enterprise tables are significantly more complex and larger. Among the three widely used benchmarks, BIRD is the most complex, with questions averaging 0.918 JOINs over 6.82 tables per database. In contrast, BEAVER questions involve 4.25 JOINs over 105.0 tables per database, making it substantially more challenging. BEAVER is still in its early stages as a benchmark. The dw dataset we used had some issues, like UNIQUE constraint violations on primary keys for which we ignored and allowed duplicates. Our long context pipeline achieves Ex Acc of 60.41% (48.64% ignoring questions with null golden queries) with gemini-1.5-pro without using the golden tables, but the entire DB tables and the column mappings, similar to the hints in BIRD benchmark. The The best 1-shot Ex Acc on BEAVER (dw + 45 additional questions), reported by (Chen et al., 2024), is 2.0 in PL/SQL using gpt-4o with relevant table schema retrieval, and 0.0 without retrieval.
The normalized generation latency $T_{g}$ increases with both larger context size and task difficulty (question and schema complexity), as it includes self-correction and retries (using exponential backoff, with a maximum of 5 retries per request). For instance, BEAVER takes significantly longer to generate on average because our long context pipeline produces context size exceeding the limits (2-million tokens for pro and 1-million tokens for flash) or results in semantically invalid (empty) outputs, in which case the model retries a fixed number of times until it finds a better answer. Consequently, for complex large datasets, like BEAVER, flash can paradoxically take as long as or even longer than pro with more retries due to the smaller context size limit.
In general, gemini-1.5-pro is better at complex reasoning tasks like NL2SQL, whereas gemini-1.5-flash is more cost-effective (lower pricing per token). We have noticed that the latest releases of the gemini models (released for public access on Sep. 24th, 2024) show significant performance improvements, and most notably, flash now performs more comparably to pro on the select benchmarks. This makes the more cost-efficient flash a highly attractive choice for production NL2SQL.
@@ 4.3. Schema linking with all database tables
We explore the impact of passing the entire table schema set collected from entire databases. Schema linking is such a critical component that we strive to achieve a high-recall in production. With a long-context model, we can include a significantly larger set of table definitions and schemas instead of limiting the input to only the most relevant top-K tables. As is standard practice, we retrieve the most relevant tables based on the NL question and the table DDL statement, using embedding similarity between the two. The top-k table retrieval is done across all the tables and DBs, mirroring the production setup illustrated in Figure 2. Alternatively, we provide all the tables from a given target DB (benchmark datasets specify to which DB the question is referring to) or from all the tables in the dataset (across all DBs), without the top-k retrieval. The goal is to ensure near perfect (higher) recall at the cost of lower precision by providing more tables, if not all, via long context .
Table 3. Table retrieval (TBR) performance and execution accuracy with the baseline pipeline. Putting all tables (all tables / DB or all tables / dataset) guarantees perfect recall
(a) BIRD dev
# tables (k) k=1 k=7 all tables / DB all tables / dataset
@@
$\bar{L}$ (tok/req) 2002.54 4627.68 7380.86 72619.96
(b) KaggleDBQA test
# tables (k) k=1 k=2 all tables / DB all tables / dataset
Ex Acc (%) 43.24 50.27 56.75 60.54
TBR Recall (%) 81.53 91.26 100 100
TBR Precision (%) 90.27 53.51 68.86 7.44
$\bar{L}$ (tok/req) 922.29 1731.64 1485.45 26087.86
The results shown in Table 3(b) demonstrate that the model does not get confused despite the presence of a large number of mostly irrelevant table definitions in the context. Thus, providing more tables, rather than solely relying on an imperfect table retrieval mechanism, can be beneficial—provided the context size allows for it. The result (for $k=1$) also re-affirms that schema linking is critical, since without all the ground truth tables in the prompt ($k=1$) the model cannot generate good quality SQL outputs. Conversely, providing all tables from the target DB ensures perfect TBR recall and results in higher Ex Acc in both BIRD dev and KaggleDBQA test . This means that TBR recall is more important than precision or accuracy as the model generation quality does not degrade much with many irrelevant tables (68.18 irrelevant tables on the average for BIRD dev and 14.875 for KaggleDBQA). The all tables / DB setting utilizes up to 13 tables per request (5 tables for KaggleDBQA), depending on the target database in BIRD dev; the average number of tables across different databases is 6.82 (2.125 for KaggleDBQA) . It is interesting to note that the execution accuracy result for all tables / dataset is slightly higher than that of all tables /DB. The difference is more pronounced in the KaggleDBQA test (almost +3.8% Ex Acc), whereas it falls within the 5% margin of error for BIRD dev (we use medium temperature for LLM generation, and Ex Acc varies slightly per run).
This suggests that the model’s output quality is at least comparable, and the inclusion of additional irrelevant schema information does not degrade generation quality.
On the one hand, including as many tables as possible (e.g., user history, all other tables within the target database) ensures high TBR recall, which long-context LLMs can accommodate. On the other hand, the latency cost increases with larger context sizes (see Fig. 5), and beyond a certain point, adding more irrelevant tables does not provide additional benefits to justify the increased cost. Furthermore, some production settings require processing hundreds of very wide tables, making it prohibitively expensive to include all tables in the context (see Section 6.3 for more discussion). Thus, accurate retrieval and filtering remains desirable, and leveraging long-context models can serve as a compensatory mechanism to offset imperfect retrieval for higher recall.
@@ 4.4. Many-shot ICL with example selection
Here, we evaluate the impact of many-shot in-context learning (ICL) using examples selected from BIRD train dataset. The examples are retrieved based on the question embedding similarity.
Table 4. Ex Acc (%) with the baseline pipeline & selected similar examples from the train datasets, $\sigma$(train), and ground-truth (+ GT)
@@
$\bar{L}$ (tok/req) 7380.86 8000.67 9927.24 13808.71 20358.30
KaggleDBQA test $\sigma$(train) 57.83 65.40 79.45 64.32 65.94
$\sigma$(train) + GT 80.50 81.62 78.91 81.08 80.50
$\bar{L}$ (tok/req) 922.29 1753.64 2550.10 4191.18 6266.45
(b) GT inserted at different positions (normalized) within the context window with selected 100 $\sigma$(train) + GT
@@
BIRD dev 77.77 78.29 78.1 78.42 78.16
KaggleDBQA test 81.62 81.62 80.00 80.00 80.08
(c) Changing the order of the examples block (100 $\sigma$(train) + GT) in the context: beginning (before the instructions), middle (before the schema details), end (after the schema details)
@@
BIRD dev 80.44 78.23 78.29
KaggleDBQA test 78.91 79.45 80.00
Table 4(c)(a) shows that the train examples did not provide much boost and instead resulted in worse performance. This means that the model’s ability to learn from similar examples from BIRD train is quite limited and it can be even distracted by them. Also, notice that the impact of distraction seems to be limited, say going from 1 similar example (11k tokens) to 100 similar examples (50k tokens). To further examine the model’s sensitivity to many irrelevant train examples, we injected 1 ground-truth example from BIRD dev in the “middle” of the example blob (the average context size is omitted as it should be almost identical to the previous table). While the model can select the ground truth example among many other irrelevant ones, the presence of those irrelevant train examples can confuse the model as opposed to just having that ground truth. The recall matters more than precision, but the model is still sensitive to bad examples (low precision) to a limited degree – note that the accuracy with 100 random examples and 1 ground truth is much higher than no examples at all.
It is also important to note that the model is robust to both the position and ordering of the ground-truth example. (Liu et al., 2024a) showed that LLMs tend to emphasize information at the beginning and end of the context window, highlighting the importance of proper ranking and ordering for retrieval-augmented generation (RAG). As in Table 4(c)(b-c), where the GT is injected at different positions within the context window, we observe that with gemini-pro-1.5 the ordering of the examples (and the schema information relocated by the relocation of the examples in (c)) has a negligible impact, and the long-context LLMs can retrieve the relevant ground truth example (also the schema information) from anywhere within its context window.
@@
(a) BIRD dev
(b) KaggleDBQA test
Figure 4. Question similarity-based example selection ($\sigma$) vs. online example generation. The baseline uses no examples; examples were retrieved from BIRD train dataset, dev dataset (excluding the GT) and synthetic generated examples. train + GT uses the retrieved train examples and the ground-truth SQL from dev.
Fig. 4 illustrates how synthetic examples improve the quality of generated SQL query and compares with other types of examples retrieved from different example pools. As is a standard practice, the example retrieval was done via question embedding similarity using gecko [^1] https://cloud.google.com/vertex-ai/generative-ai/docs/embeddings/get-text-embeddings embedding model. Note that using the ground-truth SQL queries from dev dataset provides performance ceilings ($\sigma$(dev) and $\sigma$(train + GT)), but is not allowed without an oracle (i.e., using the ground truth queries from the evaluation dataset). While dev excludes the specific GT query for a given question, it can still retrieve similar examples directly from the ground truth DB in the dev dataset. The gap between $\sigma$(dev) and $\sigma$(train + GT) indicates that providing the GT as one of the many examples is better than retrieving dev examples that are similar to the GT. $\sigma$(train) is the common practice where the relevant examples are selected from available datasets, such as the training dataset (train). In the case of BIRD dev (a), providing relevant train examples does as least as good as the baseline (no example) with a slight improvement when fewer than 200 shots are provided. However, synthetic examples ($\sigma$(synthetic) and online synthetic) provide significant uplift over $\sigma$(train). The gap between $\sigma$(synthetic) and original synthetic indicates that one can be more economical with the context size (achieving higher performance with a fewer examples) by filtering the relevant synthetic examples. In case of KaggleDBQA test (b), both training and synthetic examples boost the Ex Acc, while train performs much better than the synthetic ones. This is because BIRD train and dev datasets are sampled from different database domains, whereas KaggleDBQA train and test datasets are sampled from the same domains. It is important to note that the many-shot ICL using synthetic examples can help further improve Ex Acc, and even more so in the cross-domain case (e.g., BIRD dev) where the available training examples are less relevant.